# Graph-based Square-Root Estimation for Sparse Linear Regression

### v1

@@ Abstract
Sparse linear regression is one of the classic problems in the field of statistics, which has deep connections and high intersections with optimization, computation, and machine learning. To address the effective handling of high-dimensional data, the diversity of real noise, and the challenges in estimating standard deviation of the noise, we propose a novel and general graph-based square-root estimation (GSRE) model for sparse linear regression. Specifically, we use square-root-loss function to encourage the estimators to be independent of the unknown standard deviation of the error terms and design a sparse regularization term by using the graphical structure among predictors in a node-by-node form. Based on the predictor graphs with special structure, we highlight the generality by analyzing that the model in this paper is equivalent to several classic regression models. Theoretically, we also analyze the finite sample bounds, asymptotic normality and model selection consistency of GSRE method without relying on the standard deviation of error terms. In terms of computation, we employ the fast and efficient alternating direction method of multipliers. Finally, based on a large number of simulated and real data with various types of noise, we demonstrate the performance advantages of the proposed method in estimation, prediction and model selection.
Key words. Sparse linear regression, square-root-loss function, graphical structure among predictors, alternating direction method of multipliers, oracle property.
@@ 1 Introduction
A very active research area in statistics is the linear regression problem, whose classical model is often assumed as $y=X\beta^{*}+\sigma\epsilon$, where $y\in\mathbb{R}^{n}$ is the observed response variable, $X\in\mathbb{R}^{n\times p}$ is the design matrix, $\beta^{*}\in\mathbb{R}^{p}$ is the true coefficient to be estimated, $\sigma\geq 0$ is the noise level, and $\epsilon\in\mathbb{R}^{n}$ is an error vector of i.i.d. random variables with mean zero. Many practical applications often encounter high-dimensional small-sample data, i.e., situations with small $n$ and large $p$, which will lead to the phenomenon of rank deficiency in the design matrix $X$. Moreover, the effectiveness of traditional least squares estimation model will be greatly challenged when the predictive variables are highly correlated or the data contains non-normal distributed noise. High-dimensional statistical modeling often employs sparse penalty methods, which involves adding different penalty functions to the square loss function $\|y-X\beta\|^{2}$ for variable selection, such as $\ell_{1}$ penalty [41], $L_{p}(0<p<1)$ quasi-norm [16], smoothly clipped absolute deviation (SCAD) [14], minimax concave penalty (MCP) [53], etc. In addition, in some practical problems, such as sparse kernel learning [24], additive modeling [33] and multivariate regression [17], variable selection needs to be performed in the form of groups, which has led to the development of group penalty methods, such as group lasso [7, 51]. However, its obvious limitation lies in the non-overlapping structure, which makes it unsuitable for the practical problems where features may be encoded in multiple groups, such as many biological studies on gene interactions [29, 39].
@@
To address the aforementioned issues, we have noticed that some literature have utilized the connectivity of an undirected graph to simulate the structural information among the predictors and improve sparse regression models based on this graph. Early literature mostly used edge-by-edge method, which employs penalty functions to promote similarity in regression coefficients for correlated predictors [5, 23, 29, 49]. However, these methods did not directly utilize the neighborhood information of the graph. Additionally, when there are more edges in the graph, this approach leads to more complex penalty terms. As an improvement, [50] proposed a node-wise penalty by using a node-by-node method. Specifically, given the predictor graph $G$ and positive weights $\tau_{i}$’s, they defined the following norm
$\big{\|}\beta\big{\|}_{G,\tau}=\min\limits_{\sum_{i=1}^{p}V^{(i)}=\beta,\operatorname*{supp}(V^{(i)})\subseteq\mathcal{N}_{i}}\sum_{i=1}^{p}\tau_{i}\|V^{(i)}\|_{2},$ (1.1)
@@
where $\operatorname*{supp}(V^{(i)})$ is the support of vector $V^{(i)}\in\mathbb{R}^{p}$, $\|\cdot\|_{2}$ is the $\ell_{2}$ norm, and $\mathcal{N}_{i}$ represents the set containing the predictor $i$ and its neighbors in graph $G$. They no longer focus on the original variable $\beta$ of the model, but decompose $\beta$ into $p$ parts, and then use group sparsity idea based on predictor graph information. This design motivation comes from the explicit expression between the true coefficient and the inverse covariance matrix in analyzing least squares estimation. With the above notation, their optimization model in sparse regression incorporating graphical structure among predictors (SRIG) is given by
$\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}^{2}}{2n}+\lambda\|\beta\|_{G,\tau}\Big{\}},$ (1.2)
@@
where $\lambda$ is a regularization parameter. The SRIG method facilitates the inclusion of other predictors connected to an important predictor in a model, which is better suited to meet the demands of practical applications. Compared to penalty methods that do not use the graphical structure among predictors, SRIG method can significantly improve the performance of variable selection, estimation and prediction. The good performance of this research idea has attracted the attention of many scholars and has been extended to linear discriminant analysis [30], generalized linear regression [56, 32, 38], expectile regression [36], quantile regression [46], etc.
@@
Despite these attractive features, there still exists a theoretical and practical challenge. Because the aforementioned problems based on the smooth square loss usually need to assume that the noise follows a normal distribution, and the estimation performance depends on knowing standard deviation of the error terms in advance. The theoretical form of the regularization parameter $\lambda$, which is crucial for the model estimation, relies on the standard deviation $\sigma$. However, in high-dimensional settings, accurately estimating $\sigma$ is a non-trivial task, and the types of noise generally exhibit diversity. Fortunately, the proposal of the square-root-loss function overcomes the aforementioned limitations, which was first proposed by [35]. Based on the idea of square-root loss, [3] extended the traditional lasso problem and obtained the following square-root lasso (SRL):
$\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\|\beta\|_{1}\Big{\}},$ (1.3)
@@
where $\|\beta\|_{1}=\sum_{i=1}^{p}|\beta_{i}|$ is the $\ell_{1}$ norm of $\beta$. It has been theoretically validated that SRL can achieve near-oracle property without knowing or estimating $\sigma$ in advance. This makes the lasso-type methods much more attractive in high-dimensional scenarios. Subsequently, the square-root loss was shown to be applicable to non-normal noise under reasonable conditions [2, 11]. Furthermore, [8] extended the ideas behind the SRL for group selection. [40] proposed a nonconvex square-root-loss regression model. For more research on the square-root-loss function, one can refer to [12, 13, 37].
@@
In this paper, we carry out research on integrating the square-root-loss regression method and the penalty methods based on graphical structure among predictors. This research aims to construct a novel model for the classic high-dimensional linear regression problem, which can analyze the intrinsic information in data and alleviate the reliance on normal noise. Specifically, we propose the following graph-based square-root estimation (GSRE) model:
$\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\|\beta\|_{G,\tau}\Big{\}}.$ (1.4)
@@
To the best of our knowledge, this is the first time that square-root-loss regression incorporating graphical structure among predictors. Firstly, for several classes of predictor graphs with special structures, we verify the equivalence of the proposed GSRE method with several existing methods. Then, when the graphical structure $G$ of the predictors is known, we provide finite sample bounds, asymptotic normality and model selection consistency for the GSRE method with a $\sigma$-free tuning sequence $\lambda$. For the case where the graphical structure $G$ is unknown, we construct it by sparse estimating the covariance (or precision) matrix of the predictors [9, 18, 52]. In terms of computation, we adopt the classical alternating direction method of multiplies (ADMM) in the optimization field. Simulation studies and practical applications demonstrate that the proposed GSRE method exhibits good performance in variable selection, model estimation and prediction.
@@
Organization: The remainder of this paper is organized as follows. Relevant and rigorous theoretical results are presented in Section 2. Section 3 provides a detailed description of the classical ADMM for model (1.4). Section 4 demonstrates the numerical performance of the proposed GSRE method on both simulated and real data. Finally, we summarize this paper in Section 5.
@@
Notation: To end this section, we summarize some notations used in this paper. $[p]:=\{1,2,\ldots,p\}$. When ||$\cdot$|| ⋅ | is applied to a scalar, it denotes the absolute value, and when it acts on a set, it denotes the cardinality. For each $i\in[p]$, we denote $d_{i}$ as the number of predictors in the neighborhood ${\mathcal{N}}_{i}$, that is, $d_{i}=|{\mathcal{N}}_{i}|$. $d_{min}:=\min_{1\leq i\leq p}d_{i}$, $d_{max}:=\max_{1\leq i\leq p}d_{i}$, $\tau_{min}:=\min_{1\leq i\leq p}\tau_{i}$. For each $i\in[p]$, denote $\zeta_{i}=\|X_{{\mathcal{N}}_{i}}\|^{2}$ and $\zeta=\max_{i}\zeta_{i}$, where $\|X_{{\mathcal{N}}_{i}}\|$ is the operator norm of the matrix $X_{{\mathcal{N}}_{i}}$. For a $n\times p$ matrix $M$, $\|M\|_{\infty}$ is defined as $\max_{1\leq i\leq n}\sum_{j=1}^{p}|M_{ij}|$. For a constant $x$, we denote $(x)_{+}=\max(x,0)$. Here, we denote :=${\mathcal{I}}_{*}:=\{i:\beta_{i}^{*}\neq 0\}$, ${\mathcal{I}}_{*}^{c}:=\{i:\beta_{i}^{*}=0\}$ and $s^{*}=|{\mathcal{I}}_{*}|$ as the true nonzero coefficient set, the true zero coefficient set, and the number of true nonzero coefficients, respectively. For every $\beta\in{\mathcal{R}}^{p}$, denote $\beta_{{\mathcal{I}}_{*}}$ and $\beta_{{\mathcal{I}}_{*}^{c}}$ as the subvectors of $\beta$ with indices in ${\mathcal{I}}_{*}$ and ${\mathcal{I}}_{*}^{c}$ respectively. For any matrix $M$, $M_{{\mathcal{I}}_{*},{\mathcal{I}}_{*}}$ denote the submatrix of $M$ consisting of the entries with row and column indices in ${\mathcal{I}}_{*}$. $\beta_{min}^{*}=\min_{i\in{\mathcal{I}}_{*}}|\beta_{i}^{*}|$ denote the minimum absolute nonzero coefficient. $\operatorname*{sign}(\cdot)$ maps a positive entry to $1$, a negative entry to $-1$, and zero to zero. The symbol $\delta_{{\mathcal{C}}}(x)$ represents the indicator function over ${\mathcal{C}}$ such that $\delta_{{\mathcal{C}}}(x)=0$ if $x\in{\mathcal{C}}$ and +$\infty$+ ∞ otherwise.
@@ 2.1 Subgradient conditions
Assume there exists a vector $\beta\in\mathbb{R}^{p}$ satisfying $X\beta\neq y$. Then $\beta$ is an optimal solution of (1.4) if and only if $\beta$ can be decomposed as $\beta=\sum_{i=1}^{p}V^{(i)}$ where $V^{(i)}$’s satisfy that, for all $1\leq i\leq p$,
@@
According to Lemma 2.1, if $(\hat{V}^{(1)},\hat{V}^{(2)},\ldots,\hat{V}^{(p)})$ is a decomposition of the optimal solution of (1.4), then for each $i$, either $\hat{V}^{(i)}=\textbf{0}_{p}$ or $\operatorname*{supp}(\hat{V}^{(i)})={\mathcal{N}}_{i}$. This result is very consistent with the design idea of node-wise penalty.
@@ 2.2 Connections with some existing methods
    If the predictor graph has no edge, the problem (1.4) is the same as
    $\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\sum_{i=1}^{p}\tau_{i}|\beta_{i}|\Big{\}}.$ (2.1)
    When $\tau_{i}\equiv 1$ for each $i$, (2.1) is the classic square-root lasso [3].
@@
    If the predictor graph consists of $J$ disconnected complete subgraphs, and the nodes in the $J$ disconnected complete subgraphs are assumed to be $\{1,2,\ldots,q_{1}\},\{q_{1}+1,\ldots,q_{1}+q_{2}\},\ldots,\{q_{J-1}+1,\ldots,q_{J-1}+q_{J}\}$, then the problem (1.4) is equivalent to
    $\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\sum_{j=1}^{J}\tilde{\tau}_{j}\|\beta^{(j)}\|_{2}\Big{\}},$ (2.2)
    where $\tilde{\tau}_{j}=\min\limits_{q_{j-1}+1\leq i\leq q_{j-1}+q_{j}}\tau_{i}$ and $\beta^{(j)}=(\beta_{q_{j-1}+1},\beta_{q_{j-1}+2},\ldots,\beta_{q_{j-1}+q_{j}})^{\top}$. Obviously, this problem is exactly the group square-root lasso problem [8].
@@
    If the predictor graph is a complete graph, then the problem (1.4) has the same nonzero solutions as
    $\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\tilde{\lambda}}{n}\|\beta\|_{2}^{2}\Big{\}},$ (2.3)
@@
Lemma 2.2 shows that square-root lasso and group square-root lasso are special cases of the proposed model (1.4). This conclusion validates the broader applicability of our model, enabling it to handle predictor graphs with diverse structures.

### v2

@@ Abstract
Sparse linear regression is one of the classic problems in the field of statistics, which has deep connections and high intersections with optimization, computation, and machine learning. To address the effective handling of high-dimensional data, the diversity of real noise, and the challenges in estimating standard deviation of noise, we propose a novel and general graph-based square-root estimation (GSRE) model for sparse linear regression. Specifically, we use square-root-loss function to encourage the estimators to be independent of the unknown standard deviation of error terms and design a sparse regularization term by using the graphical structure among predictors in a node-by-node form. Based on the predictor graphs with special structure, we highlight the generality by analyzing that the model in this paper is equivalent to several classic regression models. Theoretically, we also analyze the finite sample bounds, asymptotic normality and model selection consistency of GSRE method without relying on standard deviation of error terms. In terms of computation, we employ the fast and efficient alternating direction method of multipliers. Finally, based on a large number of simulated and real data with various types of noise, we demonstrate the performance advantages of the proposed method in estimation, prediction and model selection. Supplementary materials for this article are available online.
Keywords: Sparse linear regression, square-root-loss function, graphical structure among predictors, alternating direction method of multipliers, oracle property.
@@ 1 Introduction
A very active research area in statistics is the linear regression problem, whose classical model is often assumed as $y=X\beta^{*}+\sigma\epsilon$, where $y\in\mathbb{R}^{n}$ is the observed response variable, $X\in\mathbb{R}^{n\times p}$ is the design matrix, $\beta^{*}\in\mathbb{R}^{p}$ is the true coefficient to be estimated, $\sigma\geq 0$ is the noise level, and $\epsilon\in\mathbb{R}^{n}$ is an error vector of i.i.d. random variables with mean zero. Many practical applications often encounter high-dimensional small-sample data, i.e., situations with small $n$ and large $p$, which will lead to the phenomenon of rank deficiency in the design matrix $X$. Moreover, the effectiveness of traditional least squares estimation model will be greatly challenged when the predictive variables are highly correlated or the data contains non-normal distributed noise. High-dimensional statistical modeling often employs sparse penalty methods, which involves adding different penalty functions to the square loss function $\|y-X\beta\|^{2}$ for variable selection, such as $\ell_{1}$ penalty (Tibshirani, 1996), $L_{p}(0<p<1)$ quasi-norm (Frank and Friedman, 1993), smoothly clipped absolute deviation (SCAD) (Fan and Li, 2001), minimax concave penalty (MCP) (Zhang, 2010), etc. In addition, in some practical problems, such as sparse kernel learning (Koltchinskii and Yuan, 2010), additive modeling (Meier et al., 2009) and multivariate regression (Friedman, 1991), variable selection needs to be performed in the form of groups, which has led to the development of group penalty methods, such as group lasso (Bühlmann and Geer, 2011; Yuan and Lin, 2006). However, traditional group penalty methods often require non-overlapping structure between subgroups, which makes them unsuitable for the practical problems where features may be encoded in multiple groups, such as many biological studies on gene interactions (Li and Li, 2008; Subramanian et al., 2005).
@@
To effectively handle group sparsity problems with overlapping structures between subgroups, we have noticed that some literature has utilized the connectivity of an undirected graph to simulate the structural information among the predictors and improve sparse regression models based on this graph. Early literature mostly used edge-by-edge method, which employs penalty functions to promote similarity in regression coefficients for correlated predictors (Bondell and Reich, 2008; Kim and Xing, 2009; Li and Li, 2008; Yang et al., 2012). However, these methods did not directly utilize the neighborhood information of the graph. Additionally, when there are more edges in the graph, this approach leads to more complex penalty terms. As an improvement, Yu and Liu (2016) proposed a node-wise penalty by using a node-by-node method. Specifically, given the predictor graph $G$ and positive weights $\tau_{i}$’s, they defined the following norm
$\big{\|}\beta\big{\|}_{G,\tau}=\min\limits_{\sum_{i=1}^{p}V^{(i)}=\beta,\operatorname*{supp}(V^{(i)})\subseteq\mathcal{N}_{i}}\sum_{i=1}^{p}\tau_{i}\|V^{(i)}\|_{2},$ (1)
@@
where $V^{(i)}\in\mathbb{R}^{p}$ is a partition of $\beta$, and there may be overlapping parts among the $V^{(i)}$’s, $\operatorname*{supp}(V^{(i)})$ is the support of vector $V^{(i)}\in\mathbb{R}^{p}$, $\|\cdot\|_{2}$ is the $\ell_{2}$ norm, and $\mathcal{N}_{i}$ represents the set containing the predictor $i$ and its neighbors in graph $G$. To avoid repetition, the design motivation of $\|\beta\|_{G,\tau}$ is summarized in the supplementary materials. They no longer focus on the original variable $\beta$ of the model, but decompose $\beta$ into $p$ parts, and then use group sparsity idea based on predictor graph information. With the above notation, their optimization model in sparse regression incorporating graphical structure among predictors (SRIG) is given by
$\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}^{2}}{2n}+\lambda\|\beta\|_{G,\tau}\Big{\}},$ (2)
@@
where $\lambda$ is a regularization parameter. The SRIG method facilitates the inclusion of other predictors connected to an important predictor in a model, which is better suited to meet the demands of practical applications. Compared to penalty methods that do not use the graphical structure among predictors, SRIG method can significantly improve the performance of variable selection, estimation and prediction. The good performance of this research idea has attracted the attention of many scholars and has been extended to linear discriminant analysis (Liu et al., 2019), generalized linear regression (Zhou et al., 2019; Luo and Liu, 2022; Stephenson et al., 2021), expectile regression (Pan et al., 2023), quantile regression (Wang et al., 2021), etc.
@@
Despite these attractive features, there still exists a theoretical and practical challenge. Because the aforementioned problems based on the smooth square loss usually need to assume that the noise follows a normal distribution, and the estimation performance depends on knowing standard deviation of the error terms in advance. The theoretical form of the tuning parameter $\lambda$, which is crucial for the model estimation, relies on the standard deviation $\sigma$. However, in high-dimensional settings, accurately estimating $\sigma$ is a non-trivial task, and the types of noise generally exhibit diversity. Fortunately, the proposal of the square-root-loss function overcomes the aforementioned limitations, which was first proposed by Owen (2007). Based on the idea of square-root loss, Belloni et al. (2011) extended the traditional lasso problem and obtained the following square-root lasso (SRL):
$\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\|\beta\|_{1}\Big{\}},$ (3)
@@
where $\|\beta\|_{1}=\sum_{i=1}^{p}|\beta_{i}|$ is the $\ell_{1}$ norm of $\beta$. It has been theoretically validated that SRL can achieve near-oracle property without knowing or estimating $\sigma$ in advance. This makes the lasso-type methods much more attractive in high-dimensional scenarios. Subsequently, the square-root loss was shown to be applicable to non-normal noise under reasonable conditions (Bellec et al., 2018; Cui et al., 2018). Furthermore, Bunea et al. (2013) extended the ideas behind the SRL for group selection. Tang et al. (2020) proposed a nonconvex square-root-loss regression model. For more research on the square-root-loss function, one can refer to Derumigny (2018), Ding et al. (2023) and Shen et al. (2016).
@@
In this paper, we carry out research on integrating the square-root-loss regression method and the penalty methods based on graphical structure among predictors. This research aims to construct a novel model for the classic high-dimensional linear regression problem, which can analyze the intrinsic information in data. Specifically, we propose the following graph-based square-root estimation (GSRE) model:
$\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\|\beta\|_{G,\tau}\Big{\}}.$ (4)
@@
Firstly, for several classes of predictor graphs with special structures, we verify the equivalence of the proposed GSRE method with several existing methods. Then, when the graphical structure $G$ of the predictors is known, we provide finite sample bounds, asymptotic normality and model selection consistency for the GSRE method with a $\sigma$-free tuning sequence $\lambda$. For the case where the graphical structure $G$ is unknown, we construct it by estimating sparse precision matrix of the predictors (Cai et al., 2011; Friedman et al., 2008; Yuan and Lin, 2007). In terms of computation, we adopt the classical alternating direction method of multiplies (ADMM) in the optimization field. Simulation studies and practical applications demonstrate that the proposed GSRE method exhibits good performance in variable selection, model estimation and prediction.
To the best of our knowledge, this is the first study to introduce the graphical structure among predictors into square-root-loss regression. The theoretical analysis relies on a $\sigma$-free tuning parameter, which completely eliminates the dependence on the standard deviation of the noise. Additionally, the results of this paper are also applicable to scenarios with non-Gaussian noises, which is more in line with practical needs. Furthermore, for predictors with complex correlation structures, the ingenious design of penalty terms allows for in-depth analysis and effective processing of the essential correlations. These innovations jointly expand the scope of practical applications of the research results, and provide more generalizable technical support for regression analysis in complex data environments.
@@
Organization: The remainder of this paper is organized as follows. Relevant and rigorous theoretical results are presented in Section 2. Section 3 provides a detailed description of the classical ADMM for model (4). Section 4 demonstrates the numerical performance of the proposed GSRE method on both simulated and real data. Finally, we summarize this paper in Section 5. Technical proofs are provided in the supplementary materials.
@@
Notation: To end this section, we summarize some notations used in this paper. $[p]:=\{1,2,\ldots,p\}$. When ||$\cdot$|| ⋅ | is applied to a scalar, it denotes the absolute value, and when it acts on a set, it denotes the cardinality. For each $i\in[p]$, we denote $d_{i}$ as the number of predictors in the neighborhood ${\mathcal{N}}_{i}$, that is, $d_{i}=|{\mathcal{N}}_{i}|$. $d_{min}:=\min_{1\leq i\leq p}d_{i}$, $d_{max}:=\max_{1\leq i\leq p}d_{i}$, $\tau_{min}:=\min_{1\leq i\leq p}\tau_{i}$. For each $i\in[p]$, denote $\zeta_{i}=\|X_{{\mathcal{N}}_{i}}\|^{2}$ and $\zeta=\max_{i}\zeta_{i}$, where $\|X_{{\mathcal{N}}_{i}}\|$ is the operator norm of the matrix $X_{{\mathcal{N}}_{i}}$. For a $n\times p$ matrix $M$, $\|M\|_{\infty}$ is defined as $\max_{1\leq i\leq n}\sum_{j=1}^{p}|M_{ij}|$. For a constant $x$, we denote $(x)_{+}=\max(x,0)$. Here, we denote :=${\mathcal{I}}_{*}:=\{i:\beta_{i}^{*}\neq 0\}$, ${\mathcal{I}}_{*}^{c}:=\{i:\beta_{i}^{*}=0\}$ and $s^{*}=|{\mathcal{I}}_{*}|$ as the true nonzero coefficient set, the true zero coefficient set, and the number of true nonzero coefficients, respectively. For every $\beta\in{\mathcal{R}}^{p}$, denote $\beta_{{\mathcal{I}}_{*}}$ and $\beta_{{\mathcal{I}}_{*}^{c}}$ as the subvectors of $\beta$ with indices in ${\mathcal{I}}_{*}$ and ${\mathcal{I}}_{*}^{c}$ respectively. For any matrix $M$, $M_{{\mathcal{I}}_{*},{\mathcal{I}}_{*}}$ denote the submatrix of $M$ consisting of the entries with row and column indices in ${\mathcal{I}}_{*}$. $\beta_{min}^{*}=\min_{i\in{\mathcal{I}}_{*}}|\beta_{i}^{*}|$ denote the minimum absolute nonzero coefficient. $\operatorname*{sign}(\cdot)$ maps a positive entry to 111, a negative entry to 1-1- 1, and zero to zero. The symbol $\delta_{{\mathcal{C}}}(x)$ represents the indicator function over ${\mathcal{C}}$ such that $\delta_{{\mathcal{C}}}(x)=0$ if $x\in{\mathcal{C}}$ and +$\infty$+ ∞ otherwise.
@@ 2.1 Subgradient conditions
Assume there exists a vector $\beta\in\mathbb{R}^{p}$ satisfying $X\beta\neq y$. Then $\beta$ is an optimal solution of (4) if and only if $\beta$ can be decomposed as $\beta=\sum_{i=1}^{p}V^{(i)}$ where $V^{(i)}$’s satisfy that, for all $1\leq i\leq p$,
@@
According to Lemma 2.1, if $(\hat{V}^{(1)},\hat{V}^{(2)},\ldots,\hat{V}^{(p)})$ is a decomposition of the optimal solution of (4), then for each $i$, either $\hat{V}^{(i)}=\textbf{0}_{p}$ or $\operatorname*{supp}(\hat{V}^{(i)})={\mathcal{N}}_{i}$. This result is very consistent with the design idea of node-wise penalty.
@@ 2.2 Connections with some existing methods
    If the predictor graph has no edge, the problem (4) is the same as
    $\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\sum_{i=1}^{p}\tau_{i}|\beta_{i}|\Big{\}}.$ (5)
    When $\tau_{i}\equiv 1$ for each $i$, (5) is the classic square-root lasso (Belloni et al., 2011).
@@
    If the predictor graph consists of $J$ disconnected complete subgraphs, and the nodes in the $J$ disconnected complete subgraphs are assumed to be $\{1,2,\ldots,q_{1}\},\{q_{1}+1,\ldots,q_{1}+q_{2}\},\ldots,\{q_{J-1}+1,\ldots,q_{J-1}+q_{J}\}$, then the problem (4) is equivalent to
    $\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\sum_{j=1}^{J}\tilde{\tau}_{j}\|\beta^{(j)}\|_{2}\Big{\}},$ (6)
    where $\tilde{\tau}_{j}=\min\limits_{q_{j-1}+1\leq i\leq q_{j-1}+q_{j}}\tau_{i}$ and $\beta^{(j)}=(\beta_{q_{j-1}+1},\beta_{q_{j-1}+2},\ldots,\beta_{q_{j-1}+q_{j}})^{\top}$. Obviously, this problem is exactly the group square-root lasso problem (Bunea et al., 2013).
@@
    If the predictor graph is a complete graph, then the problem (4) has the same nonzero solutions as
    $\min_{\beta\in\mathbb{R}^{p}}\Big{\{}\frac{\|y-X\beta\|_{2}}{\sqrt{n}}+\frac{\tilde{\lambda}}{n}\|\beta\|_{2}^{2}\Big{\}},$ (7)
@@
Lemma 2.2 shows that square-root lasso and group square-root lasso are special cases of the proposed model (4). This conclusion validates the broader applicability of our model, enabling it to handle predictor graphs with diverse structures.